(1 point) In this exercise we consider finding the first five coefficients in the series solution of the first order linear iPlease write neat so I can read and understand the problem.
Answer #1

no 34 - ny +4y=0 y lo)=3, y: 10)-2 let solution is y. & Conx? Now, y= Encours y (P-1) Co 29-12 - 3 D (MM) (m 20-2En Cox=> 3 en (n-1) Comento Core?+u CM 22 9 = 0 2 Now index in last two terr tror. o an Shifting to 2-2 0 - 9 1 =) 3 & n(n-1) Con-» 3 - соу - Со + Cl2 + (ә я? - 3 = 910) = Со т (1) -- C, + 2 Can + - - 2 - to) - с. » (Coe 3, С-г / Сл- (n+2 ) , Co-a Со е30»); ; com. — zo - Со + Со»?, с » ч. 4, 6, 7 * Су » Са», Cs * Су”, - 7 8. (m) е Cot tan® + Cy +9+ C6 6 + - Со [++ (-2) »?»,